Verizox Company uses a job order cost system with manufacturing overhead applied to products based on direct labor hours. At the beginning of the most recent year, the company estimated its manufacturing overhead cost at $177,510. Estimated direct labor cost was $435,540 for 18,300 hours.

Actual costs for the most recent month are summarized here:

Item Description Total Cost
Direct labor (1,830 hours) $ 43,555
Indirect costs
Indirect labor 2,420
Indirect materials 3,420
Factory rent 3,390
Factory supervision 4,770
Factory depreciation 5,750
Factory janitorial work 1,250
Factory insurance 1,760
General and administrative salaries 4,260
Selling expenses 5,330


Required:
1.
Calculate the predetermined overhead rate.

2. Calculate the amount of applied manufacturing overhead.

3. Calculate actual manufacturing overhead costs.

4. Compute over- or underapplied overhead.
Answer #1

Predetermined Overhead rate = Estimated overhead / Estimated total labor hours

= 177,510/18300

= 9.70

Applied manufacturing overhead

= 9.70*1830

= 17,751

Actual manufacturing overhead

= 2420+3420+3390+4770+5750+1250+1760

= 22,760

Underapplied overhead = 5009

(22760-17751)